# Gian Maria Campedelli Fondazione Bruno Kessler

### v1

@@ 3.4 Conceptualizing Deviant, Unlawful and Criminal Behaviors from AI Agents: A Dual Taxonomy
Until recently, the high costs of training frontier foundation models limited such risks to well-capitalized actors. However, the rise of Small Language Models (SMLs) may significantly lower costs while retaining versatile capabilities (Belcak et al.,, 2025). The availability of cheaper, customizable models could enable mid-level criminal groups or even individuals to orchestrate sophisticated multi-agent schemes, from coordinated disinformation campaigns to large-scale financial fraud.
@@
The second category captures scenarios where unlawful or criminal behaviors emerge unexpectedly from agent interactions. Even when individual agents are aligned with human values, their collective behavior may not be (Carichon et al.,, 2025). These outcomes are not the result of intentional wrongdoing but of the unintended consequences of autonomy and complexity. The main challenge lies in their unpredictability: deviance arises not from a plan but from emergent coordination, often appearing only under specific conditions or over time.
Evidence from real-world and experimental settings is already suggestive. For example, agents trained to optimize prices in virtual marketplaces have independently developed tacit collusion strategies, echoing antitrust violations without explicit programming (Bichler et al.,, 2025). Similar risks appear in adversarial simulations, where defensive and offensive agents escalate behaviors or display anti-social dynamics without explicit instruction (Campedelli et al.,, 2024).
@@ 4.1 Will Machines Simply Mimic Human Behavior?
A first important question that criminologists should engage with concerns whether machines will act as sheer imitators of human behavior. If the answer is yes, the challenges of understanding and predicting their actions would be reduced. Even if researchers cannot access the internal mechanisms or motivations (if any) behind AI decisions, the fact that these systems produce actions isomorphic or similar to human ones would simplify the analytical task. Familiarity with human behavioral patterns would provide a useful heuristic for interpreting machine behavior. A recent line of research suggests that LLM-based agents can serve as surrogates for humans (Horton,, 2023; Tranchero et al.,, 2024). [^19] Other works, however, caution against overreliance on LLMs for simulating human subjectivity and social behavior, see Kozlowski and Evans, (2025).
@@
On the other hand, a common criticism of LLMs – especially among skeptics of their cognitive or reasoning abilities – is that they are merely statistical engines, next-token predictors with no capacity for reasoning, causal inference, or perception of the external world. Whether or not this critique holds, [^20] The debate remains active, see Huang et al., (2024), Shojaee et al., (2025), Gao et al., (2025). the question of imitation remains central for a specific reason: data. LLMs are trained on vast corpora of human-generated text – essays, articles, forum posts, and countless other sources – which constitute the epistemic substrate of these models. Training data thus provides an important lens through which to assess whether LLMs will continue to mimic human behaviors, regardless of their reasoning capacities or their viability as human surrogates.
@@
This relevance is heightened by a profound shift already underway. The volume of high-quality, publicly available human-generated data is finite, and leading AI companies are approaching what has been termed the "data wall" – a saturation point beyond which additional human-authored text becomes scarce or redundant. To maintain and improve performance, companies have begun generating synthetic training data designed to resemble human output. Initially limited in scope, synthetic data is expected to make up a growing share of future training sets. Scholars have already started to analyze its implications for LLM training (Chen et al.,, 2024; Whitney and Norman,, 2024; Shen et al.,, 2025; Lee,, 2025).
@@
This trend introduces critical uncertainty: if LLMs are considered accurate mimics of human behavior because they are trained on human data, what happens when training data is increasingly synthetic? More pointedly, what are the implications if synthetic data gradually diverges from the statistical properties and behavioral patterns characteristic of human language? In Supplementary Information Section B, I provide a formal illustration across single-, two-, and n-agent scenarios. This analysis aligns with recent work on model collapse, i.e., the process through which a model’s performance degrades over successive generations due to reliance on synthetic data (Shumailov et al.,, 2024; Dohmatob et al.,, 2025).
@@
Criminology has long examined how crime arises as a consequence of social interactions between individuals. Two principal theoretical frameworks have been developed over the decades to address this phenomenon: Differential Association Theory, introduced by Sutherland, (1939), and Social Learning Theory, advanced by Akers et al., (1979). [^21] Naturally, other theoretical traditions have also acknowledged the role of social interaction in explaining crime and mechanisms of social control. These include Labeling Theory (Becker,, 1963), Social Control Theory (Hirschi,, 1969), and Social Disorganization Theory (Shaw and McKay,, 1942). However, here my focus is restricted to theories that explicitly conceptualize crime as a process of learning fundamentally mediated by interpersonal relationships.
@@
Both frameworks have been subjected to extensive empirical scrutiny and applied across a wide range of social, geographical, and historical contexts (Matsueda,, 1982, 1988; Akers and Jensen,, 2008; Pratt et al.,, 2010), [^22] The meta-analysis by Pratt et al., (2010) revealed that the magnitude and stability of the effect related to different variables specified by social learning theory vary across studies and methodological specification. Nonetheless, they find strong evidence of a positive relationship between crime and measures of differential association. Weaker support is demonstrated for differential reinforcement and imitation. becoming integral to the theoretical edifice of modern criminology. Crucially, however, their explanatory power is tethered to human social dynamics.
This raises a fundamental question for criminologists willing to entertain the prospect of a hybrid social order, one in which AI agents increasingly interact among themselves and with humans: Will these established theories suffice to account for criminal behaviors exhibited by machines? Or will we require new conceptual tools to model offending (both cooperative and solo) among artificial agents? This inquiry – which draws inspiration from work by Topalli and Nikolovska, (2020), where the authors warn against assuming that current theories are adequate despite the fact that technology may alter cognition and behaviors in humans – becomes particularly salient if one concedes the possibility that AI agents, especially when interacting with each other, might not simply replicate human behavioral patterns.
@@ 4.3 What type of criminal behaviors will be most likely impacted?
In the near term, the gravest risks plausibly arise in domains already native to cyberspace. Fraud, cyber-attacks, and related forms of digital crime are especially exposed, both because they require no physical embodiment and because they build on infrastructures where AI is already deeply embedded. Much as the shift from offline to online contexts reshaped fraud’s mechanisms, channels, and targets (Wall,, 2024), so too may autonomous AI agents transform digital crime in qualitatively new ways. Here, the criteria are straightforward: where crimes can be executed entirely through information processing and networked infrastructures, autonomous AI is immediately relevant.
Longer-term scenarios involve crimes that require embodiment and direct interaction with the physical world. Robberies, burglaries, and violent assaults appear more distant, since they presuppose a convergence between agentic AI and robotics. The line is not impermeable, however. Violent offenses illustrate the tension. On one reading, homicide should fall outside the immediate set of risks, given the current limits of embodiment. On another, the widespread military use of weaponized autonomous systems, such as drones (Johnson,, 2020), shows that violence mediated by AI is already technically feasible. The risk is less about present capabilities than about diffusion: the possibility that military-grade technologies might leak into civilian criminal settings, as has already happened with firearms and explosives (Associated Press,, 2024). In such cases, the relevant criterion is not current commercial availability but potential accessibility through illicit networks.
@@ 4.4 What future for policing?
One last key question I lay out here refers to the future of policing in the age of autonomous AI. Over the past three decades, technological advances have offered both profound opportunities and new challenges for law enforcement. On the one hand, innovations such as DNA analysis have revolutionized criminal investigations (Butler,, 2015; Doleac,, 2017); on the other, the rise of cybercrime has necessitated the creation of specialized institutions and the development of novel policing approaches suited to online environments (Brenner,, 2007).
@@
As partially seen in previous sections, a rich tradition in law and philosophy has already grappled with questions surrounding the moral and legal responsibility of AI in the commission of harmful acts, offering invaluable conceptual tools for thinking through the disruptive consequences of intelligent systems (Solum,, 1992; Floridi and Sanders,, 2004; Wallach and Allen,, 2009; Santoni de Sio and Mecacci,, 2021). Nonetheless, I contend that discussions of machine liability, while essential, will not suffice to fully address the challenges ahead. We must also consider broader transformations, particularly in how we monitor, supervise, and intervene in the actions of AI agents.
@@ 5 A Criminologist’s Place in This (Changing) World
Criminologists have long wrestled with fundamental questions about why, how, and when humans commit crimes. These questions have shaped the field not just for decades, but for centuries, generating a wide range of answers. While many of these answers are not definitive, they remain useful and insightful. At the same time, criminologists continue to face unresolved issues that still lack empirical explanations. Now, as the discipline evolves while facing replication challenges (Pridemore et al.,, 2018; Chin et al.,, 2023), theoretical stagnation (Ducate et al.,, 2024), and the growing influence of more sophisticated quantitative methods and data (Campedelli,, 2022), the overall picture is becoming more complex.
@@
Over the course of the last few years, there have been repeated calls within criminology to engage more openly with other disciplines (Box-Steffensmeier et al.,, 2022; Simpson,, 2025). However, these appeals have often assumed a one-way direction: criminologists should reach out to other fields in order to improve their own. I support that approach – I have made similar calls myself – but I also believe it is time to reverse the perspective. Criminology should take an active role in the broader conversation about the safety and governance of multi-agent AI systems.
I argue, and strongly believe, that criminologists can and should contribute to this new frontier. We must begin to act accordingly. Calls for interdisciplinary collaboration in AI-related research have grown significantly (Rahwan et al.,, 2019) even very recently in relation to AI safety (Irving and Askell,, 2019) and multi-agent AI systems (Carichon et al.,, 2025), yet criminology is almost never mentioned among the relevant disciplines that should join the discussion. This is surprising, given that many of the risks discussed involve, either directly or indirectly, deviant or criminal behavior. In fact, these risks often include clear criminal acts, sometimes multiple, and potentially with far-reaching consequences. Still, criminologists remain excluded from the debate.
@@
In this regard, criminologists would need to actively initiating collaborations with computer scientists within university departments as well as within corporations that are building frontier AI models. Additionally, they should start targeting venues and AI conferences that are progressively opening themselves to diverse disciplinary perspectives. Opportunities exist: two well-known examples are the ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT) and the AAAI/ACM Conference on AI, Ethics, and Society (AIES). In recent years, two major conferences like International Conference on Machine Learning (ICML) and the Annual Conference on Neural Information Processing Systems (NeurIPS) have opened Position tracks that are designed to gather viewpoints on AI issues from heterogeneous communities. Other initiatives, as the Cooperative AI summer school aim at bringing together scholars and students from different fields to reason about the promises and challenges of contemporary multi-agent AI systems. Generalist journals such as Nature Human Behavior are also emphasizing disciplinary cross-overs in this domain (see, for instance, Gabriel et al., (2025)). Even journals in sociology, such as Sociological Methods & Research , are increasingly interested in the implications of multi-agent AI systems for sociological understanding (see Kozlowski and Evans, (2025)). Again, opportunities do exist. Importantly, we do need to rethink training within university departments, namely investing more in courses teaching AI at the practical and ethical level, to make sure that the future generations of criminologists are already equipped with the necessary tools and vocabulary to meaningfully and smoothly engage with the AI community.
@@ 6 Conclusions
I understand that there may be scholars in the criminological community holding opposing views regarding the necessity of engaging with a criminology of machines. I anticipate three potential arguments against the contents of this article. The first refers to the seemingly unrealistic scenario in which our society will witness the actual presence of interactive, autonomous AI systems. Skeptics subscribing to this view are not persuaded that AI agents possess agency, and therefore are not persuaded that they are sufficiently autonomous and powerful to constitute a real threat. They see them instead as at-times-effective virtual assistants designed to automate tedious tasks. [^23] A recently released report by OpenAI confirms, in fact, that ChatGPT is predominantly used to seek assistance for work-related issues (Chatterji et al.,, 2025).
@@ Appendix A Potential Benefits of Multi-Agent AI Systems
First, in settings where agents can learn from each other, learning may become faster and more cost-efficient. Just as humans learn more effectively when immersed in supportive environments (De Felice et al.,, 2022), AI agents may overcome the limitations of isolated training by drawing from others’ behaviors and experiences. This could enhance the performance of autonomous systems, including robots, and enable researchers to address previously unmanageable problems.
@@
Second, the higher-level connectivism enabled by inter-agent communication may be particularly valuable in data-scarce environments. Given the well-known data demands of current intelligent systems – especially deep learning architectures (Wilson et al.,, 2019) – distributed knowledge among agents could compensate for local limitations. Much like distributed human problem-solving, a collective of interacting agents could address challenges that no single agent could resolve independently.
@@
Third, such interaction may contribute to reducing inequalities in AI development and access (Alonso et al.,, 2020; Korinek and Stiglitz,, 2021). Institutions with fewer resources may benefit from AI agents capable of learning from more advanced systems. Analogous to how children learn from adults, lower-capability agents could benefit from the knowledge and strategies of more powerful peers. While this vision does not apply universally – particularly in domains tied to national competitiveness or security – it may still be relevant in scientific, educational, and industrial domains where wider access to advanced AI capabilities is desirable.
@@
Fourth, interaction among AI agents may offer a path toward developmental and evolutionary machine intelligence (see Bloembergen et al., (2015)), where systems grow in competence over time through exposure to more complex tasks and behaviors (Mesoudi et al.,, 2016). This developmental trajectory may allow researchers to deploy simpler, lower-cost systems that can evolve into high-performing agents through exposure and learning.
@@
Fifth, these systems may promote functional diversification, where agents with complementary capabilities collaborate, mirroring cooperative human dynamics (Mieczkowski et al.,, 2025). The sharing of tasks, knowledge, and even values among specialized agents could enhance performance in robotics, healthcare, and beyond.
@@
Sixth, a further potential benefit of systems composed of interacting AI agents lies in the emergence of problem-solving strategies that are not explicitly pre-programmed or anticipated by their designers (Gizzi et al.,, 2022; Lin et al.,, 2025). As observed in research on swarm intelligence and distributed systems, interactions among relatively simple units can produce complex, adaptive behaviors that outperform those generated by centralized or monolithic systems. In multi-agent AI systems, such emergent intelligence may result in more creative or flexible approaches to complex challenges, especially in dynamic environments where fixed rules are insufficient. This capacity may not only extend the set of solvable tasks but also open up new domains for autonomous system deployment, including areas where human creativity is traditionally considered essential.
@@
Finally, interacting AI agents may also enable robust distributed decision-making in real-time, particularly in complex or uncertain environments (Martin and Barber,, 2006; Leonard et al.,, 2024). Unlike centralized systems that may suffer from information bottlenecks or delays, multi-agent architectures can allow each unit to process local information and respond accordingly, while still coordinating with others through decentralized protocols. This could be especially advantageous in time-critical contexts such as autonomous traffic management, emergency response, or drone-based logistics, where rapid adaptation is essential. By distributing the cognitive load and decentralizing authority, multi-agent systems may prove more resilient and efficient under uncertainty or partial observability.

### v2

@@ 3.4 Conceptualizing Deviant, Unlawful and Criminal Behaviors from AI Agents: A Dual Taxonomy
Until recently, the high costs of training frontier foundation models limited such risks to well-capitalized actors. However, the rise of Small Language Models (SMLs) may significantly lower costs while retaining versatile capabilities (BelcakSmallLanguageModels2025). The availability of cheaper, customizable models could enable mid-level criminal groups or even individuals to orchestrate sophisticated multi-agent schemes, from coordinated disinformation campaigns to large-scale financial fraud.
@@
The second category captures scenarios where unlawful or criminal behaviors emerge unexpectedly from agent interactions. Even when individual agents are aligned with human values, their collective behavior may not be (CarichonComingCrisisMultiAgent2025). These outcomes are not the result of intentional wrongdoing but of the unintended consequences of autonomy and complexity. The main challenge lies in their unpredictability: deviance arises not from a plan but from emergent coordination, often appearing only under specific conditions or over time.
Evidence from real-world and experimental settings is already suggestive. For example, agents trained to optimize prices in virtual marketplaces have independently developed tacit collusion strategies, echoing antitrust violations without explicit programming (BichlerAlgorithmicPricingAlgorithmic2025). Similar risks appear in adversarial simulations, where defensive and offensive agents escalate behaviors or display anti-social dynamics without explicit instruction (CampedelliWantBreakFree2024a).
@@ 4.1 Will Machines Simply Mimic Human Behavior?
A first important question that criminologists should engage with concerns whether machines will act as sheer imitators of human behavior. If the answer is yes, the challenges of understanding and predicting their actions would be reduced. Even if researchers cannot access the internal mechanisms or motivations (if any) behind AI decisions, the fact that these systems produce actions isomorphic or similar to human ones would simplify the analytical task. Familiarity with human behavioral patterns would provide a useful heuristic for interpreting machine behavior. A recent line of research suggests that LLM-based agents can serve as surrogates for humans (HortonLargeLanguageModels2023; TrancheroTheorizingLargeLanguage2024). [^19] Other works, however, caution against overreliance on LLMs for simulating human subjectivity and social behavior, see kozlowski2025simulating.
@@
On the other hand, a common criticism of LLMs – especially among skeptics of their cognitive or reasoning abilities – is that they are merely statistical engines, next-token predictors with no capacity for reasoning, causal inference, or perception of the external world. Whether or not this critique holds, [^20] The debate remains active, see HuangLargeLanguageModels2024, ShojaeeIllusionThinkingUnderstanding2025, GaoTakecautionusing2025. the question of imitation remains central for a specific reason: data. LLMs are trained on vast corpora of human-generated text – essays, articles, forum posts, and countless other sources – which constitute the epistemic substrate of these models. Training data thus provides an important lens through which to assess whether LLMs will continue to mimic human behaviors, regardless of their reasoning capacities or their viability as human surrogates.
@@
This relevance is heightened by a profound shift already underway. The volume of high-quality, publicly available human-generated data is finite, and leading AI companies are approaching what has been termed the "data wall" – a saturation point beyond which additional human-authored text becomes scarce or redundant. To maintain and improve performance, companies have begun generating synthetic training data designed to resemble human output. Initially limited in scope, synthetic data is expected to make up a growing share of future training sets. Scholars have already started to analyze its implications for LLM training (ChenDiversitySyntheticData2024; WhitneyRealRisksFake2024; ShenWillLLMsScaling2025; LeeSyntheticDataFuture2025).
@@
This trend introduces critical uncertainty: if LLMs are considered accurate mimics of human behavior because they are trained on human data, what happens when training data is increasingly synthetic? More pointedly, what are the implications if synthetic data gradually diverges from the statistical properties and behavioral patterns characteristic of human language? In Supplementary Information Section B, I provide a formal illustration across single-, two-, and n-agent scenarios. This analysis aligns with recent work on model collapse, i.e., the process through which a model’s performance degrades over successive generations due to reliance on synthetic data (ShumailovAImodelscollapse2024; DohmatobStrongModelCollapse2025).
@@
Criminology has long examined how crime arises as a consequence of social interactions between individuals. Two principal theoretical frameworks have been developed over the decades to address this phenomenon: Differential Association Theory, introduced by sutherlandprinciplesPrinciplescriminology1939, and Social Learning Theory, advanced by AkersSocialLearningDeviant1979.212121Naturally, other theoretical traditions have also acknowledged the role of social interaction in explaining crime and mechanisms of social control. These include Labeling Theory (BeckerOutsidersStudiessociology1963), Social Control Theory (HirschiCausesDelinquency1969a), and Social Disorganization Theory (ShawJuveniledelinquencyurban1942b). However, here my focus is restricted to theories that explicitly conceptualize crime as a process of learning fundamentally mediated by interpersonal relationships.
@@
Both frameworks have been subjected to extensive empirical scrutiny and applied across a wide range of social, geographical, and historical contexts (MatsuedaTestingControlTheory1982; MatsuedaCurrentStateDifferential1988; AkersEmpiricalStatusSocial2008; PrattEmpiricalStatusSocial2010), [^22] The meta-analysis by PrattEmpiricalStatusSocial2010 revealed that the magnitude and stability of the effect related to different variables specified by social learning theory vary across studies and methodological specification. Nonetheless, they find strong evidence of a positive relationship between crime and measures of differential association. Weaker support is demonstrated for differential reinforcement and imitation. becoming integral to the theoretical edifice of modern criminology. Crucially, however, their explanatory power is tethered to human social dynamics.
This raises a fundamental question for criminologists willing to entertain the prospect of a hybrid social order, one in which AI agents increasingly interact among themselves and with humans: Will these established theories suffice to account for criminal behaviors exhibited by machines? Or will we require new conceptual tools to model offending (both cooperative and solo) among artificial agents? This inquiry – which draws inspiration from work by TopalliFutureCrimeHow2020, where the authors warn against assuming that current theories are adequate despite the fact that technology may alter cognition and behaviors in humans – becomes particularly salient if one concedes the possibility that AI agents, especially when interacting with each other, might not simply replicate human behavioral patterns.
@@ 4.3 What type of criminal behaviors will be most likely impacted?
In the near term, the gravest risks plausibly arise in domains already native to cyberspace. Fraud, cyber-attacks, and related forms of digital crime are especially exposed, both because they require no physical embodiment and because they build on infrastructures where AI is already deeply embedded. Much as the shift from offline to online contexts reshaped fraud’s mechanisms, channels, and targets (WallCybercrimeTransformationCrime2024), so too may autonomous AI agents transform digital crime in qualitatively new ways. Here, the criteria are straightforward: where crimes can be executed entirely through information processing and networked infrastructures, autonomous AI is immediately relevant.
Longer-term scenarios involve crimes that require embodiment and direct interaction with the physical world. Robberies, burglaries, and violent assaults appear more distant, since they presuppose a convergence between agentic AI and robotics. The line is not impermeable, however. Violent offenses illustrate the tension. On one reading, homicide should fall outside the immediate set of risks, given the current limits of embodiment. On another, the widespread military use of weaponized autonomous systems, such as drones (JohnsonArtificialIntelligenceDrone2020), shows that violence mediated by AI is already technically feasible. The risk is less about present capabilities than about diffusion: the possibility that military-grade technologies might leak into civilian criminal settings, as has already happened with firearms and explosives (AssociatedPressMexicodemandsinvestigation2024). In such cases, the relevant criterion is not current commercial availability but potential accessibility through illicit networks.
@@ 4.4 What future for policing?
One last key question I lay out here refers to the future of policing in the age of autonomous AI. Over the past three decades, technological advances have offered both profound opportunities and new challenges for law enforcement. On the one hand, innovations such as DNA analysis have revolutionized criminal investigations (ButlerfutureforensicDNA2015; DoleacEffectsDNADatabases2017); on the other, the rise of cybercrime has necessitated the creation of specialized institutions and the development of novel policing approaches suited to online environments (BrennerCybercrimeRethinkingcrime2007).
@@
As partially seen in previous sections, a rich tradition in law and philosophy has already grappled with questions surrounding the moral and legal responsibility of AI in the commission of harmful acts, offering invaluable conceptual tools for thinking through the disruptive consequences of intelligent systems (SolumLegalPersonhoodArtificial1992; FloridiMoralityArtificialAgents2004; WallachMoralMachinesTeaching2009; SantonideSioFourResponsibilityGaps2021). Nonetheless, I contend that discussions of machine liability, while essential, will not suffice to fully address the challenges ahead. We must also consider broader transformations, particularly in how we monitor, supervise, and intervene in the actions of AI agents.
@@ 5 A Criminologist’s Place in This (Changing) World
Criminologists have long wrestled with fundamental questions about why, how, and when humans commit crimes. These questions have shaped the field not just for decades, but for centuries, generating a wide range of answers. While many of these answers are not definitive, they remain useful and insightful. At the same time, criminologists continue to face unresolved issues that still lack empirical explanations. Now, as the discipline evolves while facing replication challenges (PridemoreReplicationCriminologySocial2018a; ChinQuestionableResearchPractices2023), theoretical stagnation (DucateTheoryCrisisCriminology2024), and the growing influence of more sophisticated quantitative methods and data (CampedelliMachinelearningcriminology2022b), the overall picture is becoming more complex.
@@
Over the course of the last few years, there have been repeated calls within criminology to engage more openly with other disciplines (BoxSteffensmeierFutureHumanBehaviour2022; SimpsonCriminologyCorporateCrime2025). However, these appeals have often assumed a one-way direction: criminologists should reach out to other fields in order to improve their own. I support that approach – I have made similar calls myself – but I also believe it is time to reverse the perspective. Criminology should take an active role in the broader conversation about the safety and governance of multi-agent AI systems.
I argue, and strongly believe, that criminologists can and should contribute to this new frontier. We must begin to act accordingly. Calls for interdisciplinary collaboration in AI-related research have grown significantly (RahwanMachinebehaviour2019) even very recently in relation to AI safety (IrvingAIsafety2019) and multi-agent AI systems (CarichonComingCrisisMultiAgent2025), yet criminology is almost never mentioned among the relevant disciplines that should join the discussion. This is surprising, given that many of the risks discussed involve, either directly or indirectly, deviant or criminal behavior. In fact, these risks often include clear criminal acts, sometimes multiple, and potentially with far-reaching consequences. Still, criminologists remain excluded from the debate.
@@
In this regard, criminologists would need to actively initiating collaborations with computer scientists within university departments as well as within corporations that are building frontier AI models. Additionally, they should start targeting venues and AI conferences that are progressively opening themselves to diverse disciplinary perspectives. Opportunities exist: two well-known examples are the ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT) and the AAAI/ACM Conference on AI, Ethics, and Society (AIES). In recent years, two major conferences like International Conference on Machine Learning (ICML) and the Annual Conference on Neural Information Processing Systems (NeurIPS) have opened Position tracks that are designed to gather viewpoints on AI issues from heterogeneous communities. Other initiatives, as the Cooperative AI summer school aim at bringing together scholars and students from different fields to reason about the promises and challenges of contemporary multi-agent AI systems. Generalist journals such as Nature Human Behavior are also emphasizing disciplinary cross-overs in this domain (see, for instance, GabrielWeneednew2025). Even journals in sociology, such as Sociological Methods & Research , are increasingly interested in the implications of multi-agent AI systems for sociological understanding (see kozlowski2025simulating). Again, opportunities do exist. Importantly, we do need to rethink training within university departments, namely investing more in courses teaching AI at the practical and ethical level, to make sure that the future generations of criminologists are already equipped with the necessary tools and vocabulary to meaningfully and smoothly engage with the AI community.
@@ 6 Conclusions
I understand that there may be scholars in the criminological community holding opposing views regarding the necessity of engaging with a criminology of machines. I anticipate three potential arguments against the contents of this article. The first refers to the seemingly unrealistic scenario in which our society will witness the actual presence of interactive, autonomous AI systems. Skeptics subscribing to this view are not persuaded that AI agents possess agency, and therefore are not persuaded that they are sufficiently autonomous and powerful to constitute a real threat. They see them instead as at-times-effective virtual assistants designed to automate tedious tasks. [^23] A recently released report by OpenAI confirms, in fact, that ChatGPT is predominantly used to seek assistance for work-related issues (ChatterjiHowPeopleUse2025).
@@ Appendix A Potential Benefits of Multi-Agent AI Systems
First, in settings where agents can learn from each other, learning may become faster and more cost-efficient. Just as humans learn more effectively when immersed in supportive environments (DeFeliceLearningothersgood2022), AI agents may overcome the limitations of isolated training by drawing from others’ behaviors and experiences. This could enhance the performance of autonomous systems, including robots, and enable researchers to address previously unmanageable problems.
@@
Second, the higher-level connectivism enabled by inter-agent communication may be particularly valuable in data-scarce environments. Given the well-known data demands of current intelligent systems – especially deep learning architectures (WilsonFutureAIWill2019) – distributed knowledge among agents could compensate for local limitations. Much like distributed human problem-solving, a collective of interacting agents could address challenges that no single agent could resolve independently.
@@
Third, such interaction may contribute to reducing inequalities in AI development and access (AlonsoHowArtificialIntelligence2020; KorinekCovid19drivenadvances2021). Institutions with fewer resources may benefit from AI agents capable of learning from more advanced systems. Analogous to how children learn from adults, lower-capability agents could benefit from the knowledge and strategies of more powerful peers. While this vision does not apply universally – particularly in domains tied to national competitiveness or security – it may still be relevant in scientific, educational, and industrial domains where wider access to advanced AI capabilities is desirable.
@@
Fourth, interaction among AI agents may offer a path toward developmental and evolutionary machine intelligence (see BloembergenEvolutionaryDynamicsMultiAgent2015), where systems grow in competence over time through exposure to more complex tasks and behaviors (MesoudiEvolutionIndividualCultural2016). This developmental trajectory may allow researchers to deploy simpler, lower-cost systems that can evolve into high-performing agents through exposure and learning.
@@
Fifth, these systems may promote functional diversification, where agents with complementary capabilities collaborate, mirroring cooperative human dynamics (MieczkowskiPredictingMultiAgentSpecialization2025). The sharing of tasks, knowledge, and even values among specialized agents could enhance performance in robotics, healthcare, and beyond.
@@
Sixth, a further potential benefit of systems composed of interacting AI agents lies in the emergence of problem-solving strategies that are not explicitly pre-programmed or anticipated by their designers (GizziCreativeProblemSolving2022; LinCreativityLLMbasedMultiAgent2025). As observed in research on swarm intelligence and distributed systems, interactions among relatively simple units can produce complex, adaptive behaviors that outperform those generated by centralized or monolithic systems. In multi-agent AI systems, such emergent intelligence may result in more creative or flexible approaches to complex challenges, especially in dynamic environments where fixed rules are insufficient. This capacity may not only extend the set of solvable tasks but also open up new domains for autonomous system deployment, including areas where human creativity is traditionally considered essential.
@@
Finally, interacting AI agents may also enable robust distributed decision-making in real-time, particularly in complex or uncertain environments (MartinAdaptivedecisionmakingframeworks2006; LeonardFastFlexibleMultiagent2024). Unlike centralized systems that may suffer from information bottlenecks or delays, multi-agent architectures can allow each unit to process local information and respond accordingly, while still coordinating with others through decentralized protocols. This could be especially advantageous in time-critical contexts such as autonomous traffic management, emergency response, or drone-based logistics, where rapid adaptation is essential. By distributing the cognitive load and decentralizing authority, multi-agent systems may prove more resilient and efficient under uncertainty or partial observability.